Five-sixths of a number is -60. What is the number? I have set it up like this:
(5\6)N = -60 I get here and do not know how to proceed.

Found 2 solutions by user_dude2008, stanbon:
Answer by user_dude2008(1825)

(5/6)N=-60
6*(5/6)N=-60*6
5N=-360
(5N)/5=-360/5
N=-72 <--------- Answer

Answer by stanbon(48508)

(5\6)N = -60 I get here and do not know how to proceed.
---
Multiply both sides by (6/5) to get:
N = -60(6/5)
N = -12*6
N = -72
==============
Cheers,
stan H.